# RLHFuse: Efficient RLHF Training for Large Language Models with Inter- and Intra-Stage Fusion

### v2

@@ Abstract
Reinforcement Learning from Human Feedback (RLHF) enhances the alignment between LLMs and human preference. The workflow of RLHF typically involves several models and tasks in a series of distinct stages. Existing RLHF training systems view each task as the smallest execution unit and thus overlooking the opportunities for subtask-level optimizations. Due to the intrinsic nature of RLHF training, i.e., the data skewness in the generation stage, and the pipeline bubbles in the training stage, existing RLHF systems suffer from low GPU utilization in production deployments.
RLHFuse breaks the traditional view of RLHF workflow as a composition of individual tasks, splitting each task into finer-grained subtasks, and performs stage fusion to improve GPU utilization. RLHFuse contains two key ideas. First, for generation and inference tasks, RLHFuse splits them into sample-level subtasks, enabling efficient inter-stage fusion to mitigate the original generation bottleneck dominated by long-tailed samples. Second, for training tasks, RLHFuse breaks them into subtasks of micro-batches. By leveraging the intuition that pipeline execution can be essentially complemented by another pipeline, RLHFuse performs intra-stage fusion to concurrently execute these subtasks in the training stage with a fused pipeline schedule, resulting in fewer pipeline bubbles. In addition, RLHFuse incorporates a series of system optimizations tailored for each stage of RLHF, making it efficient and scalable for our internal product usage. We evaluate RLHFuse on various popular LLMs and the results show that RLHFuse increases the training throughput by up to 3.7×3.7$\times$3.7 ×, compared to existing state-of-the-art systems.
@@ 1 Introduction
Figure 1: RLHF iteration breakdown of existing frameworks.
The rise of Large Language Models (LLMs) marks a revolutionary leap in generative AI. From search engines [1] to personal assistants [2], and from operating systems [3] to embodied AI [4], LLMs are redefining what is possible, opening up new frontiers and possibilities. Despite the impressive capabilities of LLMs, concerns around their safety, robustness, and reliability have been at the forefront since their inception. Many studies [5, 6, 7] have shown that LLMs often display unintended behaviors, such as fabricating facts, generating biased or toxic content, or failing to adhere to user instructions. To address these issues, Reinforcement Learning from Human Feedback (RLHF) [8] has been introduced, aiming to align LLMs with human intent after pre-training. With a concrete and well-defined reward target, RLHF enhances the ability of LLMs in multiple domains, including following instructions, suppressing hallucinations, and improving overall safety. It is known to be the key driving force behind the evolution of GPT-3 [9] into ChatGPT [2], and almost all the leading LLMs [10, 11, 12] incorporate RLHF to boost their model performance.
Unlike pre-training, RLHF training is significantly more complex in terms of system design, since it involves four different LLMs, three stages, and six tasks in total. As shown in Figure 1, each iteration begins with the Actor model, which generates responses to a batch of input prompts (known as the generation stage). These responses are then passed to the Reference, Critic, and Reward model for processing, during the inference stage. Finally, the Actor and Critic models compute loss and update their parameters, completing the training stage. The data and weight dependencies create synchronization boundaries between stages.
@@
To enable efficient training of RLHF, existing frameworks have proposed a series of optimizations, such as selecting tailored parallel strategies for each task [13] and optimizing the data exchange during stage transitions [14]. However, these techniques fundamentally treat the RLHF workflow as a simple composition of individual tasks, failing to delve into the inherent characteristics and structure inside the tasks, thereby missing significant optimization opportunities. In production deployments, we have observed two issues within current frameworks that severely impact training efficiency, as indicated by the grey area in Figure 1.
@@
The first issue arises from the data skewness in the generation stage, where the output length of generated samples follows a long-tailed distribution. This phenomenon is broadly observed during LLM generation (§2.2) and this data diversity is critical to the robustness of RL training. As a result, the generation time is dominated by the longest sample as shown in Figure 1. During the latter part of the generation stage, the GPUs remain idle or are engaged in processing only a small number of long-tailed samples. Due to data dependencies, the inference tasks cannot start in advance, leading to low GPU utilization. Making the problem even worse, as the maximum output length of LLMs continues to grow [15], this long-tail phenomenon will become more significant.
@@
The second issue stems from the pipeline bubbles in the training stage. With the explosive growth in the size of LLMs, higher pipeline parallelism (PP) is needed to scale training. When LLMs reach hundreds of billions of parameters, it is quite common for PP size to reach a few dozen [16, 17]. However, the proportion of pipeline bubbles increases with PP size, significantly reducing training efficiency. In RLHF, there are two training tasks, each may evolve a specific LLM with hundreds of billions of parameters, which amplifies this inefficiency. Despite that various approaches have been proposed to reduce pipeline bubbles when training individual models [18, 19, 20, 21], the remaining bubbles still pose a significant challenge under synchronous training semantics.
@@
To this end, we propose RLHFuse, the first work to employ subtask-level optimizations for RLHF training. By breaking down the task into finer-grained subtasks, RLHFuse opens up a new design space to fuse the stage execution. This is orthogonal to the state-of-the-art RLHF framework ReaLHF [13] that optimizes the parallelism strategy for each task. RLHFuse exploits the optimizations opportunities at the subtask level with two new techniques, i.e., data-aware inter-stage fusion and model-aware intra-stage fusion. These two techniques leverage the unique system characteristics of RLHF training to address data skewness and pipeline bubbles, respectively.
@@
For data skewness, we split the generation and inference tasks into sample-level subtasks and then the dependency granularity between the generation and inference stage can be refined from stage-level to sample-level. This is due to the fact that the computation of the two stages is essentially independent for different samples. To leverage this insight, we design a lightweight and efficient migration algorithm to automatically detect when the generation stage enters the inefficient long-tail processing phase and migrates the remaining long-tailed samples to a limited number of GPUs. The freed-up resources are used to launch the inference tasks. This approach enables inter-stage fusion and dynamically adjusts the migration timing in each iteration according to the workload, thereby maximizing overall efficiency.
@@
For pipeline bubbles, we break the training task into subtasks of micro-batches. Pipeline bubbles are essentially caused by dependencies between subtasks of the same model. Fortunately, in the RLHF training stage, there are two independent models, and their subtasks are naturally independent as well. Based on this observation, we employ a fused pipeline schedule to enable intra-stage fusion, which executes the training tasks on the same set of GPUs, effectively filling each other’s pipeline bubbles. We design a lightweight pipeline schedule generation algorithm that can produce near-optimal schedules for Actor and Critic models of any size and parallel configurations.
@@ 2.1 Background
RLHF models. Typically, RLHF training involves four LLMs in total: Actor, Ref, Critic, and RW. The Actor model, which is the primary model to be trained and the final product after RLHF, serves as the agent in the RL semantics. Given a prompt, each token it generates autoregressively is considered an action, and the goal of the PPO algorithm is to guide it to produce actions that align with the reward target. The Reference model (Ref) is initialized from the original Actor model, but is frozen (i.e., weights do not update) during training. It provides Kullback-Leibler (KL) divergence regularization to ensure the Actor model does not deviate excessively from its original version during training or generate nonsensical outputs. The Reward model (RW) is trained on human-labeled preference data to score each sample generated by the Actor model, which guides the Actor model to generate responses that align with human preferences. It is also frozen during the RLHF training. The Critic model, initialized from the RW model, serves as the value model to evaluate the actions taken by the Actor. It provides finer-grained feedback on action level to guide Actor towards better decisions.
RLHF workflow. Figure 1 illustrates the workflow of one training iteration in RLHF, which contains three main stages: Generation, Inference, and Training.
Generation stage. The Actor model generates responses for the prompts in the current batch. This process involves a prefill phase and a decoding phase. The prefill phase processes the prompt, which is composed of a sequence of tokens, to generate the first output token. Following this, the decoding phase sequentially generates subsequent tokens in multiple steps autoregressively (i.e., each decoding step generates a new token based on tokens generated in previous steps until reaching a termination token). Ultimately, each prompt and its corresponding response form one training sample for this iteration, referred to as a rollout or trajectory in the context of RL. To increase sample efficiency, it is common to generate hundreds or even thousands of samples in each iteration.
@@
Training stage. In the PPO algorithm, unlike traditional LLM training, all the samples are divided into several mini-batches and the model weights need to update after training on each mini-batch. Concretely, for each mini-batch, the Actor and Critic models first perform a forward pass, then calculate the loss using the results from the inference stage and perform a backward pass to update the parameters. After the Actor completes training on all mini-batches, it uses the up-to-date parameters for the generation in the next iteration.
LLM parallelization. Guided by the scaling law [23], the size of LLMs has increased exponentially. Various parallelism methods are proposed to scale training, including data parallelism, tensor parallelism, and pipeline parallelism.
Data Parallelism (DP). DP replicates the model weights and distribute the data among the replicas to execute in parallel.At the end of each iteration, all DP groups perform all-reduce communication on the gradients before updating parameters.
Tensor Parallelism (TP). TP partitions computationally intensive operators, mainly matrix multiplications, across multiple GPUs to accelerate computation. It requires significant communication, so it is typically used within a node to leverage the high-bandwidth NVLINK connections.
Pipeline Parallelism (PP). PP organizes LLM layers into stages, each running on a separate device or node. It partitions the input batch into multiple “micro-batches” to form pipeline execution and accumulates the gradients of the entire input batch. PP features point-to-point communication and incurs much less communication overhead than TP.
Since the above parallelism methods have their own advantages, all three parallel strategies are used together to scale training in practice, such as Megatron-LM [16] and MegaScale [17] in pre-training [16] and ReaLHF [13] in RLHF training. We use an integer triple $(dp,pp,tp)$ to denote the degrees of DP, PP, and TP in 3D-parallelism.
@@ 2.2 Problems in RLHF Training
Generation Stage: data skewness. One major problem in the generation stage is that the response length of generated samples exhibits a long-tailed distribution, meaning that each batch will contain a few samples that are significantly longer than the others. Due to data dependencies, the inference tasks cannot start until the generation task finishes. Consequently, even only a few long-tailed samples force the inference tasks to wait. This leads to significantly low GPU utilization in the generation stage because the decoding phase of LLM generation is memory-bound, requiring a large batch size (typically a few hundred) to maintain high GPU utilization.
@@
Figure 2 (right) presents the iteration time breakdown under different maximum output length settings. The time spent generating long-tailed samples (dark blue bar) accounts for more than half of the total generation time. The situation gets worse as the maximum generation length increases, which leads to a substantial rise in iteration time. It is worth noting that this change only affects the long-tailed samples ($<1\%$), but the results indicate that this very small portion of long-tail samples can have a significant impact on the overall system performance. As user applications increasingly demand lengthy outputs [15], the inclusion of samples with long response lengths in RLHF training becomes more common and important, making the long-tail phenomenon an urgent problem in real deployment.
@@
Figure 3: The timeline of 1F1B and interleaved 1F1B pipeline schedule with 4 pipeline stages and 4 micro-batches. Interleaved 1F1B deploys two model chunks on each pipeline stage in this example.
Interleaved 1F1B scheduling [16] is proposed to mitigate the pipeline bubbles. As shown in Figure 3 (bottom), it divides the LLM into more fine-grained model chunks, with each stage hosting $K(K>1)$ model chunks, reducing the pipeline bubble percentage to $\frac{N-1}{N-1+KM}$ However, it not only introduces a $K$-fold communication overhead but also exacerbates the imbalances across PP stages. As a result, $K$ is typically set to 2 in practice, leaving room for further optimizations.
There are other works that aim to completely eliminate pipeline bubbles; however, they either violate the synchronous training semantics [25] or rely on specific, often unpractical, assumptions [21]. As a result, pipeline bubbles still pose a significant problem in LLM training, especially when the model size keeps scaling.
Summary. Essentially, the above issues are caused by the inherent characteristics of RLHF, which are difficult to eliminate through task-level optimizations. Existing systems treat the task as the smallest execution unit and overlook its internal structure, thereby missing huge optimization opportunities.
@@ 3 RLHFuse Overview
Workflow. With specific algorithm, model, and cluster configurations as input, RLHFuse first configures efficient parallel and deployment strategies following the approach in ReaLHF [13]. For each task within the RLHF workflow, RLHFuse will assign it a tailored parallel strategy $P_{i}=(dp,pp,tp)$ to maximize the GPU utilization. To deploy the task onto the physical cluster, we introduce the concept of device mesh, which is the unit for executing an individual task. A device mesh $D=(N,M)$ is defined as a two-dimensional grid of GPUs, covering $N$ nodes where each is equipped with $M$ devices. RLHFuse will designate a device mesh $D_{i}$ to each task which can be executed with parallel strategy $P_{i}$. At runtime, RLHFuse will launch each task with its corresponding parallel strategy on the designated device mesh following the workflow dependency and handle the weight redistribution and data transmission between different tasks to ensure synchronous training semantics. Meanwhile, RLHFuse utilizes data-aware inter-stage fusion to fuse the generation and inference tasks and model-aware intra-stage fusion to fuse the training tasks.
Inter-stage fusion. As depicted in the left part of Figure 4, during the generation stage, there are multiple model instances, each managing complete model weights and part of the samples. RLHFuse actively monitors each generation instance and triggers migration when the number of remaining samples is below a particular threshold. The remaining samples will be migrated to some dedicated generation instances which are designated to handle long-tailed samples. Subsequently, the resources of the original generation instances are repurposed to launch inference tasks, thereby allowing for the overlap of inference tasks with the generation of long-tailed samples and fusing the execution of the two stages.
Intra-stage fusion. As shown in the right part of Figure 4, in the training stage, RLHFuse leverages the insight that pipeline execution can be complemented by another pipeline and generates a fused pipeline schedule based on the model size and parallel strategies of Actor and Critic model. It utilizes a light-weight schedule generation algorithm to minimize the pipeline bubbles and activation memory usage, thus improving the training efficiency.
Figure 4: RLHFuse architecture.

### v3

# Optimizing RLHF Training for Large Language Models with Stage Fusion

@@ Abstract
We present RLHFuse, an efficient training system with stage fusion for Reinforcement Learning from Human Feedback (RLHF). Due to the intrinsic nature of RLHF training, i.e., the data skewness in the generation stage and the pipeline bubbles in the training stage, existing RLHF systems suffer from low GPU utilization. RLHFuse breaks the traditional view of RLHF workflow as a composition of individual tasks, splitting each task into finer-grained subtasks, and performing stage fusion to improve GPU utilization. RLHFuse contains two key ideas. First, for generation and inference tasks, RLHFuse splits them into sample-level subtasks, enabling efficient inter-stage fusion to overlap the execution of generation and inference stages, thus mitigating the original generation bottleneck dominated by long-tailed samples. Second, for training tasks, RLHFuse breaks them into subtasks of micro-batches and performs intra-stage fusion to concurrently execute these subtasks in the training stage with a fused pipeline schedule, effectively mitigating the pipeline bubbles. The experiments show that RLHFuse increases the training throughput by up to 3.7×3.7$\times$3.7 ×, compared to existing systems.
@@ 1 Introduction
The rise of Large Language Models (LLMs) marks a revolutionary leap in generative AI. Despite the impressive capabilities of LLMs, many studies [1, 2, 3] have shown that LLMs often display unintended behaviors. To address these issues, Reinforcement Learning from Human Feedback (RLHF) [4] has been introduced, aiming to align LLMs with human intent after pre-training. With a concrete and well-defined reward target, RLHF enhances the ability of LLMs in multiple domains. Figure 1 shows the typical RLHF training workflow, which involves six different tasks divided into three stages. The Actor model first generates samples with the input prompts and passes them to the Reference, Critic, and Reward models for inference, finally the loss is calculated to train both the Actor and Critic models, while the other two models remain frozen. There are barriers between stages due to the data and weight dependencies.
Figure 1: RLHF workflow [4] and the problems in the generation and training stages of existing RLHF training frameworks.
@@
To enable efficient training of RLHF, existing frameworks  [5, 6, 7] have proposed a series of optimizations, such as selecting tailored parallel strategies for each task and optimizing the data exchange during stage transitions. However, these techniques fundamentally treat the RLHF workflow as a simple composition of individual tasks, failing to delve into the inherent characteristics and structure inside the tasks, thereby missing significant optimization opportunities. In production deployments, we have observed two issues within current frameworks that severely impact training efficiency, as indicated by the grey area in Figure 1.
@@
The first issue arises from the data skewness in the generation stage, where the output length of generated samples follows a long-tailed distribution. This phenomenon is broadly observed during LLM generation (§2.2) and this data diversity is critical to the robustness of RL training. On the one hand, during the latter part of the generation stage, only a small number of long-tailed samples remain processing, but a large batch size is required for generation to improve GPU utilization (§2.1). On the other hand, later tasks cannot start in advance due to data dependencies. As a result, the generation time is dominated by the longest sample and the GPU utilization is low, as shown in the upper right of Figure 1. Although numerous optimization techniques [8, 9, 10, 11, 12] are proposed to improve the generation speed and throughput, they cannot alter the inherent long-tail phenomenon and thus fail to solve the problem. Making the problem even worse, as the maximum output length of LLMs continues to grow [13, 14], this long-tail phenomenon will become more significant.
@@
The second issue stems from the pipeline bubbles in the training stage. With the explosive growth in the size of LLMs, higher pipeline parallelism (PP) size is needed to scale training. When LLMs reach hundreds of billions of parameters, it is quite common for PP size to reach a few dozen [15, 16]. However, the proportion of pipeline bubbles increases with PP size, significantly reducing training efficiency. In RLHF, there are two training tasks, each may evolve a specific LLM with hundreds of billions of parameters, which amplifies this inefficiency. Despite that various approaches have been proposed to reduce pipeline bubbles when training individual models [17, 18, 19, 20], the remaining bubbles still pose a significant challenge under synchronous training semantics.
@@
To this end, we propose RLHFuse, which tackles the long-tailed generation issue and pipeline bubbles by employing subtask-level optimizations. By breaking down the task into finer-grained subtasks, RLHFuse opens up new design space to fuse the stage execution. This is orthogonal to current RLHF frameworks[5, 6, 7] that optimizes on the task-level to choose the optimal parallel strategies for each task or reduce the task switching overhead. RLHFuse focuses on how to execute the RLHF workflow efficiently with stage fusion after the parallel strategies are chosen. RLHFuse exploits the opportunities at the subtask level and proposes two new techniques, i.e., data-aware inter-stage fusion and model-aware intra-stage fusion. These two techniques leverage the unique system characteristics of RLHF training to fuse the execution across different tasks, thus effectively addressing the data skewness and pipeline bubbles, respectively.
@@
For data skewness, we split the generation and inference tasks into sample-level subtasks and then the dependency granularity between the generation and inference stage can be refined from stage-level to sample-level. This is due to the fact that the computation of the two stages is essentially independent for different samples. To leverage this insight, we design a lightweight and efficient migration algorithm to automatically detect when the generation stage enters the inefficient long-tail processing phase and migrate the remaining long-tailed samples to a limited number of GPUs. The freed-up resources are used to launch the inference tasks. This approach enables inter-stage fusion and dynamically adjusts the migration timing in each iteration according to the workload, thereby maximizing overall efficiency.
@@
For pipeline bubbles, we break the training task into subtasks of micro-batches. Pipeline bubbles are essentially caused by dependencies between subtasks of the same model. Fortunately, in the RLHF training stage, there are two independent models, and their subtasks are mutually independent. Based on this observation, we employ a fused pipeline schedule to enable intra-stage fusion, which executes the training tasks on the same set of GPUs with bidirectional pipelines, effectively filling each other’s pipeline bubbles. We design a lightweight schedule generation algorithm that can produce near-optimal schedules for Actor and Critic models of any size and parallel configurations.
@@ 2.1 Background
RLHF models. Typically, RLHF training involves four LLMs in total: Actor, Ref, Critic, and RW. The Actor model, which is the primary model to be trained and the final product after RLHF, serves as the agent in the RL semantics. Given a prompt, each token it generates autoregressively is considered an action, and the goal of the PPO algorithm is to guide it to produce actions that align with the reward target. The Reference model (Ref) is initialized from the original Actor model, but is frozen (i.e., weights do not update) during training. It provides Kullback-Leibler (KL) divergence regularization to ensure the Actor model does not deviate excessively from its original version during training or generate nonsensical outputs. The Reward model (RW) is trained on human-labeled preference data to score each sample generated by the Actor model, which guides the Actor model to generate responses that align with human preferences. It is also frozen during the RLHF training. The Critic model, initialized from the RW model, serves as the value model to evaluate the actions taken by the Actor. It provides finer-grained feedback on the action level to guide Actor towards better decisions.
RLHF workflow. The workflow of one RLHF training iteration contains three main stages:
Generation stage. The Actor model generates responses for the prompts in the current batch. This process involves a prefill phase and a decoding phase. The prefill phase processes the prompt to generate the first output token. Then the decoding phase sequentially generates subsequent tokens in multiple steps autoregressively (i.e., each decoding step generates a new token based on tokens generated in previous steps until reaching a termination token). Ultimately, each prompt and its corresponding response form one training sample for this iteration, referred to as a rollout or trajectory in the context of RL. To increase sample efficiency, it is common to generate many samples in one iteration.
@@
Training stage. In the PPO algorithm, unlike traditional LLM training, all the samples are divided into several mini-batches and the model weights need to be updated after training on each mini-batch. Concretely, for each mini-batch, the Actor and Critic models first perform a forward pass, then calculate the loss using the results from the inference stage and performing a backward pass to update the parameters. After the Actor completes training on all mini-batches, it uses the up-to-date parameters for the generation in the next iteration.
LLM parallelization. Various parallelism methods are proposed to scale training, guided by the scaling law [22]. Data Parallelism (DP) replicates the model weights and distributes the data among the replicas to execute in parallel. It requires gradient synchronization across the replicas after each iteration. Tensor Parallelism (TP) assigns individual operators over multiple GPUs, with each executing part of the computation in parallel. It requires significant communication, so it is typically used within a node to leverage the high-bandwidth NVLINK connections. Pipeline Parallelism (PP) organizes LLM layers into stages, each running on a separate device or node. It partitions the input batch into multiple “micro-batches” to form pipeline execution and accumulate the gradients of the entire input batch. Each parallel method has its own advantages, so in practice all three parallel strategies are used together to scale training, such as Megatron-LM [15] and MegaScale [16] in pre-training [15] as well as ReaLHF [6] and HybridFlow [5] in RLHF training.
@@ 2.2 Problems in RLHF Training
Generation Stage: data skewness. One major problem in the generation stage is that the response length of generated samples exhibits a long-tailed distribution, meaning that a few samples are significantly longer than the others. Due to data dependencies, the inference tasks cannot start until the generation task finishes. Consequently, even only a few long-tailed samples force the inference tasks to wait. This leads to significantly low GPU utilization in the generation stage because the decoding phase of LLM generation is memory-bandwidth-bound, requiring a large batch size (typically a few hundred) to maintain high GPU utilization.
@@
Figure 2 (right) presents the iteration time breakdown under different maximum output length settings. The time spent generating long-tailed samples (dark blue bar) accounts for more than half of the total generation time. The situation gets worse as the maximum generation length increases, which leads to a substantial rise in iteration time. It is worth noting that this change only affects the long-tailed samples ($<1\%$), but the results indicate that this very small portion of long-tailed samples can have a significant impact on the overall system performance. As user applications increasingly demand lengthy outputs [13, 14, 24], the inclusion of samples with long response lengths in RLHF training becomes more common and important, making the long-tail phenomenon an urgent problem in real deployment.
@@
Figure 3: The timeline of 1F1B [18] and interleaved 1F1B [15] pipeline schedule with 4 pipeline stages and 4 micro-batches.
Interleaved 1F1B scheduling [15] is proposed to mitigate the pipeline bubbles. As shown in Figure 3 (bottom), it divides the LLM into more fine-grained model chunks, with each stage hosting $K(K>1)$ model chunks, reducing the pipeline bubble percentage to $\frac{N-1}{N-1+KM}$ However, it not only introduces a $K$-fold communication overhead but also exacerbates the imbalances across PP stages. As a result, $K$ is typically set to a small constant in practice, leaving room for further optimizations.
Other works [18, 20] aim to eliminate pipeline bubbles; however, they either violate the synchronous training semantics [18] or rely on specific assumptions [20]. As a result, pipeline bubbles still pose a significant problem in LLM training, especially when the model size keeps scaling.
Summary. Essentially, the above issues are caused by the inherent characteristics of RLHF, which are difficult to eliminate through task-level optimizations. Existing systems treat the task as the smallest execution unit and overlook its internal structure, thereby missing a huge optimization opportunities.
Figure 4: RLHFuse architecture.
@@ 3 RLHFuse Overview
Workflow. With specific algorithm, model, and cluster configurations as input, RLHFuse first configures efficient parallel and deployment strategies following the approach in ReaLHF [6] and HybridFlow [5]. For each task within the RLHF workflow, RLHFuse assigns it a tailored parallel strategy to maximize GPU utilization. At runtime, RLHFuse launches each task with its corresponding parallel strategy on the designated device mesh following the workflow dependency and handle the weight redistribution and data transmission between different tasks to ensure synchronous training semantics. Meanwhile, RLHFuse utilizes data-aware inter-stage fusion to fuse the generation and inference tasks and model-aware intra-stage fusion to fuse the training tasks.
Inter-stage fusion. As depicted in the left part of Figure 4, during the generation stage, there are multiple model instances, each managing complete model weights and part of the samples. RLHFuse actively monitors each generation instance and triggers migration when the number of remaining samples is below a particular threshold. The remaining samples are migrated to some dedicated generation instances which are designated to handle long-tailed samples. Subsequently, the resources of the original generation instances are repurposed to launch inference tasks, thereby allowing for the overlap of inference tasks with the generation of long-tailed samples and fusing the execution of the two stages.
Intra-stage fusion. As shown in the right part of Figure 4, in the training stage, RLHFuse leverages the insight that another inverse pipeline can complement current pipeline execution. It utilizes a lightweight algorithm to generate a fused pipeline schedule based on the model size and parallel strategies of the Actor and Critic model, which minimizes the pipeline bubbles and activation memory usage.